### v1

@@ Abstract
Keywords:
Internet of Things Stream Processing Active Inference Autoscaling Markov Decision Processes Reinforcement Learning
@@ 1 Introduction
However, Edge and CC environments introduce new challenges [1]: They rely on resource-constrained computing hardware, and thus break with traditional Cloud-based autoscaling. In Cloud systems, autoscaling mechanisms elastically respond to increased user demand by allocating more resources to a service or replicating it. This is infeasible at the Edge or in the CC, where computing resources are strictly limited [9]. Especially when resources are scarce, applications require a more flexible scaling behavior that uses a wider range of adaptations – hence, operating in multiple elasticity dimensions [3]. On the one hand, this protects the service execution and promises higher requirements fulfillment – captured through a set of Service Level Objectives (SLOs). On the other hand, this increases the complexity for choosing optimal scaling actions. What is needed, hence, are lightweight multi-dimensional scaling mechanisms that optimize the service execution without obstructing existing workloads.
To fill this gap, we propose an agent-based autoscaling approach tailored for Edge and CC systems, which adjusts processing services in multiple elasticity dimensions. Our approach employs decentralized local agents that (1) observe the service execution and their SLO fulfillment without centralized control; thus, we can monitor the resource allocation per service or the application throughput. If SLOs are violated, the agents attempt to restore the desired state by (2) adjusting the service execution; the exact scaling policy is learned by the agent according to environmental feedback. Notably, our approach allows scaling policies tailored to the individual services, where one service could, for example, scale down its machine learning (ML) model, while another service claim the remaining resources. This allows building composite and customizable scaling policies, which go further than existing approaches [13].
@@ 2.2 Related Work
Most notably, competing approaches and traditional autoscaling methods [17, 18] struggle in environments, where multi-dimensional adaptation is crucial for maintaining service quality. In particular, RL-based systems still find no productive use by large providers despite over a decade of research. We argue that AIF-based agents offer a promising alternative by enabling adaptive, efficient, and decentralized control suited to the dynamic and uncertain nature of complex multi-tier distributed systems. Hence, our focus is on further establishing AIF for service adaptation on resource-restricted devices.
Among existing works, Sedlak et al. [14] explore how AIF can be used to optimize SLO fulfillment across various use cases. Danilenka et al. [2] demonstrate how AIF agents can adapt to dynamic and heterogenous environments. Pujol et al. [12] adopt a representation using Partially Observable Markov Decision Processes (POMDPs) within a multi-agent system. Lapkovskis et al. [8] presented a comparison of AIF agents with other methods for a CC application that complements our work with a different set of algorithms. Vyas et al. [16] provide an adaptation mechanism for active sensing on Edge devices, which dynamically adjust perception by changing the camera orientation.
@@ 3.1.1 Processing Services
quality int [128, 320] $\checkmark$ (step$=32$) $\geq 288$
@@
quality int [300, 1000] $\checkmark$ (step$=100$) $\geq 900$
@@ 3.3.2 Deep Active Inference (DACI)
While traditional AIF implementations using graphical models provide excellent interpretability, they are challenging to scale, limiting their practicality in real-world applications with high-dimensional input spaces. DACI addresses scalability in the CC by exploiting the abundant availability of hardware acceleration (e.g., TPUs, NPUs) for Artificial Neural Networks (ANNs). This work presents preliminary results on Deep Active Inference for the CC based on the work by Fountas et al [5]. The basic idea is to learn non-linear transforms with ANNs to map the high-dimensional input space into a compressed latent representation of the environment. Then, the agent may operate efficiently in complex, multi-dimensional elasticity spaces that would be intractable for discrete graphical models. Analogous to any active inference agent, it 1) samples the environment and calibrates its internal generative model to explain sensory observations and 2) performs actions to reduce the uncertainty about the environment. We slightly simplify and adjust the original objective in [5] to compute the variational free energy for each time-step t as:
$\mathcal{F}^{\prime}_{t}=$ $\mathbb{E}_{Q_{\phi_{s}}(s_{t})}\left[\log P_{\theta_{o}}(o_{t}|s_{t})\right]$
@@
We have omitted the KL divergence involving the habitual network and included regularization terms to reduce the variance between service fulfillments and encourage resource utilization. Our intuition is to regularize the term $-\mathbb{E}[\log P(o_{\tau}|\pi)]$ that represents preferences about future observations by encoding the specification for balanced services and resource efficiency to have a higher prior probability. Lastly, we encode the objective to prioritize actions that reduce uncertainty about the environment by minimizing the EFE with an expression that is equivalent to that in the original work [5].
@@ 3.3.3 Deep Q Networks (DQN)
The DQN [10] agent approximates Q-values for state-action pairs in the multi-dimensional autoscaling problem. We implement separate DQN models for each processing service (CV and QR), allowing service-specific policy learning while maintaining coordinated resource allocation when competing for limited resources. The models are trained jointly within the shared processing environment to benefit from the pre-training in the simulated LGBN environment, resulting in the DQN agent achieving stable performance from the initial deployment phase.

### v2

@@ 1 Introduction
However, Edge and CC environments introduce new challenges [1]: They rely on resource-constrained computing hardware, and thus break with traditional Cloud-based autoscaling. In Cloud systems, autoscaling mechanisms elastically respond to increased user demand by allocating more resources to a service or replicating it. This is infeasible at the Edge or in the CC, where computing resources are strictly limited [manaouil_kubernetes_2020]. Especially when resources are scarce, applications require a more flexible scaling behavior that uses a wider range of adaptations – hence, operating in multiple elasticity dimensions [2]. On the one hand, this protects the service execution and promises higher requirements fulfillment – captured through a set of Service Level Objectives (SLOs). On the other hand, this increases the complexity for choosing optimal scaling actions. What is needed, hence, are lightweight multi-dimensional scaling mechanisms that optimize the service execution without obstructing existing workloads.
To fill this gap, we propose an agent-based autoscaling approach tailored for Edge and CC systems, which adjusts processing services in multiple elasticity dimensions. Our approach employs decentralized local agents that (1) observe the service execution and their SLO fulfillment without centralized control; thus, we can monitor the resource allocation per service or the application throughput. If SLOs are violated, the agents attempt to restore the desired state by (2) adjusting the service execution; the exact scaling policy is learned by the agent according to environmental feedback. Notably, our approach allows scaling policies tailored to the individual services, where one service could, for example, scale down its machine learning (ML) model, while another service claim the remaining resources. This allows building composite and customizable scaling policies, which go further than existing approaches [sedlak_towards_2025_short].
@@ 2.2 Related Work
Most notably, competing approaches and traditional autoscaling methods [12, zhao_tiny_2022_short] struggle in environments, where multi-dimensional adaptation is crucial for maintaining service quality. In particular, RL-based systems still find no productive use by large providers despite over a decade of research. We argue that AIF-based agents offer a promising alternative by enabling adaptive, efficient, and decentralized control suited to the dynamic and uncertain nature of complex multi-tier distributed systems. Hence, our focus is on further establishing AIF for service adaptation on resource-restricted devices.
Among existing works, Sedlak et al. [9] explore how AIF can be used to optimize SLO fulfillment across various use cases. Danilenka et al. [danilenka_adaptive_2024_short] demonstrate how AIF agents can adapt to dynamic and heterogenous environments. Pujol et al. [8] adopt a representation using Partially Observable Markov Decision Processes (POMDPs) within a multi-agent system. Lapkovskis et al. [6] presented a comparison of AIF agents with other methods for a CC application that complements our work with a different set of algorithms. Vyas et al. [11] provide an adaptation mechanism for active sensing on Edge devices, which dynamically adjust perception by changing the camera orientation.
@@ 3.1.1 Processing Services
quality int [128, 320] $\checkmark$ (step=32=32) $\geq 288$
@@
quality int [300, 1000] $\checkmark$ (step=100=100) $\geq 900$
@@ 3.3.2 Deep Active Inference (DACI)
While traditional AIF implementations using graphical models provide excellent interpretability, they are challenging to scale, limiting their practicality in real-world applications with high-dimensional input spaces. DACI addresses scalability in the CC by exploiting the abundant availability of hardware acceleration (e.g., TPUs, NPUs) for Artificial Neural Networks (ANNs). This work presents preliminary results on Deep Active Inference for the CC based on the work by Fountas et al [mcdaci]. The basic idea is to learn non-linear transforms with ANNs to map the high-dimensional input space into a compressed latent representation of the environment. Then, the agent may operate efficiently in complex, multi-dimensional elasticity spaces that would be intractable for discrete graphical models. Analogous to any active inference agent, it 1) samples the environment and calibrates its internal generative model to explain sensory observations and 2) performs actions to reduce the uncertainty about the environment. We slightly simplify and adjust the original objective in [mcdaci] to compute the variational free energy for each time-step t as:
$\mathcal{F}^{\prime}_{t}=\$$\mathbb{E}_{Q_{\phi_{s}}(s_{t})}\left[\log P_{\theta_{o}}(o_{t}|s_{t})\right]$
@@
We have omitted the KL divergence involving the habitual network and included regularization terms to reduce the variance between service fulfillments and encourage resource utilization. Our intuition is to regularize the term $\mathbb{E}[\log P(o_{\tau}|\pi)]$ that represents preferences about future observations by encoding the specification for balanced services and resource efficiency to have a higher prior probability. Lastly, we encode the objective to prioritize actions that reduce uncertainty about the environment by minimizing the EFE with an expression that is equivalent to that in the original work [mcdaci].
@@ 3.3.3 Deep Q Networks (DQN)
The DQN [dqn] agent approximates Q-values for state-action pairs in the multi-dimensional autoscaling problem. We implement separate DQN models for each processing service (CV and QR), allowing service-specific policy learning while maintaining coordinated resource allocation when competing for limited resources. The models are trained jointly within the shared processing environment to benefit from the pre-training in the simulated LGBN environment, resulting in the DQN agent achieving stable performance from the initial deployment phase.
@@ 5 Conclusion & Future Work
Acknowledgment
This work is partially supported by CNS2023-144359 and the European Union NextGenerationEU/PRTR under MICIU/AEI/10.13039/501100011033, and partially by the European Union under (TEADAL, 101070186).